Are the following statements true?
 40 persons : 200 persons = ₹ 15 : ₹ 75 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

40 persons : 200 persons =₹15 :₹75
40 / 200 = 1 / 5
15 / 75 = 1 / 5
Hence, True

Thanks